HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  laser spectroscopy on Be-like Kr, Page 4 of 6  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  43   Fri Feb 28 00:18:16 2025 TaminaGeneralKalibration UV Monitor im LaserlaborDer UV Monitor steht in der Reflexion des
PBSC auf dem Lasertisch im Laserlabor. 
  
  42   Thu Feb 27 23:01:26 2025 Danyal, Konstantin, Sebastian CalibrationDelaytime measurementYesterday, we forgot to take the transit time
of the PMT's into account when we set up
the delay times of the detectors PMT middle,
  
  41   Thu Feb 27 21:23:02 2025 JensDAQPython Data ChannelsPython variables and names in DAQ
Because the Name of variables in DAQ doesnt
match the the real variables, here is List
  
  40   Thu Feb 27 20:21:31 2025 SebastianGeneral10 MHz Ref Clock (Butis)10 MHz Referenz Clock von BUTIS mit AFG als
externl clock verbunden.
Wird benötigt das die Bunch Kavität (DDS
  
  39   Thu Feb 27 16:30:57 2025 Konstantin und RodolfoDAQHF SignalWe have exchanged the HF-Signal in the DAQ
Electronic: from Dummy-HF to the real HF
from accelerator. 
  
  38   Thu Feb 27 11:19:55 2025 DanyalAcceleratorpictures of the ion beam from 26.02.2025 Hier sind noch ein paar Bilder vom Ionenstrahl
, wie er gestern Abend war.
 Schottky_pick-up_Bild.jpgSchottky_resonator_Bild.jpggebunchter_Strahl_im_ESR.jpg 
  37   Thu Feb 27 10:52:32 2025 RodolfoDetectorsTaking PMTs into operation[B]PMT Nord[/B] the power supply (CH0) is
giving an alarm already by an applied voltage
of -1707 V (operation voltage is -2450 V).
 20250227_105411.jpg20250227_143518.jpg 
  36   Thu Feb 27 09:59:08 2025 KonstantinCalibrationHV-DividerThe readout of the HV divider runs on a RaspberryPi
inside the HV-cave of the electron cooler.
Starting from 2025-02-27 09:42, the divider
  
  35   Thu Feb 27 09:38:04 2025 Konstantin & DanyalDetectorsLaufzeit MessungenGestern Abend haben wir die Signal Laufzeiten
im gemessen:
  
  34   Wed Feb 26 18:53:52 2025 DanyalAcceleratorESR settingsDate: 26.02.2025 @  18:35

E-cooler:  HV = 157366 +/- 1 V  (HV divider
  
  33   Tue Feb 25 14:51:52 2025 DanyalGeneralExperiment AcceptanceToday, 25 FEB 2025 at 9:00 AM, our experiment
was reviewed carefully by the safety committee.
  
  32   Tue Feb 25 13:12:01 2025 SebastianDAQTeilungsverhältnis Ohmlabs HVS250Hochspannung Maßstabsfaktoren

-140 kV >>> 248514
 Bericht_PTB_verbesserter_Teiler_2018.pdfMaßstabsfaktoren.JPG 
  31   Thu Feb 20 13:51:04 2025 Denise, SebatianDAQDAQ Channelliste[TABLE border="1"]
TDC Channels |
|- 
  
  30   Thu Feb 20 13:37:30 2025 Denise, SebastianDAQCFD Schwellen DAQSchwellen fuer den CFD N842 (lascool)
Lasereinstellungen dabei waren: 0.5 Skalenteile
-> 115ps Pulse
  
  29   Thu Feb 20 00:45:00 2025 DanyalGeneralDoppler width of M1 transitionAccording to Paul Indelicato

transition 3P1 -> 3P0  in 78Kr32+ : 
  
  28   Mon Feb 17 13:49:20 2025 DeniseGeneralFrequenzbreiten IonensystemBerechnung der Frequenzbreiten des Lasers
im Ionensystem 
Werte für die Frequezbreiten des Lasers aus
  
  27   Fri Feb 14 16:41:54 2025 Sebastian GeneralVerkabelung PD ESRVerkabelung von den Photodioden im ESR. Sind
auch gelabelt.
 IMG_8368.jpeg 
  26   Fri Feb 14 16:24:22 2025 DeniseGeneralStrahlprofil UV im ESR Erstes Bild ist das Strahlptofil in Höhe des
Nord 
towers. 
 IMG_8360.jpeg796a74f2-ec76-4396-82e5-9d3209b36612.jpeg0bbd839a-f6cf-4fa4-a5a4-71d0c84044fa.mov 
  25   Thu Feb 13 12:56:01 2025 DanyalGeneralESR access - monday (whole day) 8-18 hr    
  24   Thu Feb 13 12:55:13 2025 DanyalGeneralESR access - tomorrow (Fri) 10:00 - 12:00    
ELOG V3.1.5-fc6679b